OpenStudy (sonyalee77):

Two tenths of a percent of what is 15? To solve this problem, use the basic percent equation. 0.2 x n = 15 A.True B. False

OpenStudy (anonymous):

percent literally means per 100, or 1/100. two tenths of 1% is the same as \[\frac{2}{10}\times \frac{1}{100}\]

OpenStudy (jack1):

took the words right out of my mouth @hea

OpenStudy (anonymous):

.02x = 15

OpenStudy (sonyalee77):

so its true

OpenStudy (anonymous):

No. To find x, first multiply both sides by 100

OpenStudy (sonyalee77):

oh thanx you guys you've helped me alot

Directrix (directrix):

Isn't two tenths of one percent equal to .002 or two thousandths.

OpenStudy (anonymous):

\[\frac{2}{1000} \times n = 15\] should be what you get @Directrix is a boss

OpenStudy (anonymous):

-"Isn't two tenths of one percent equal to .002 or two thousandths." DOH!!!! I *HATE* when I don't read the question properly!!! Yes, @Directrix is right! My example was only for 2%, not for 2/10 of a percemt

OpenStudy (anonymous):

Sot the expression should be .002x = 15 And to solve for x, first multiply both sides by 1000

Directrix (directrix):

So, I guess the answer is FALSE. These decimals and percentages, especially when mixed, are tricky. So, no worries about an errant decimal point. Shake if off and move on to the next problem, I say.